# Argumentative Debates for Transparent Bias Detection [Technical Report]

### v1

@@ Abstract
As the use of AI in society grows, addressing emerging biases is essential to prevent systematic disadvantages against specific groups. Several fairness notions/algorithmic methods have been proposed, but, with very few exceptions, these tend to ignore transparency. Instead, interpretability and explainability are core requirements for algorithmic fairness, even more so than for other algorithmic solutions, given the human-oriented nature of fairness. In this paper, we contribute a novel transparent method for bias detection to empower debates about the presence of bias against individuals, based on the values of protected features for the individuals and others in their neighbourhoods. Our method builds upon techniques from (formal and computational) argumentation, whereby debates result from arguing about biases within and across neighbourhoods. We evaluate our approach experimentally and demonstrate its strengths in performance against an argumentative baseline.
@@ 1 Introduction
Figure 1: A QBAF generated by our approach for the COMPAS dataset (with protected feature $X_{p}\!=\!race$ and protected group g=g\!= African-American) for two neighbourhoods of size $K=10,100$ (critical question arguments with zero strength omitted). Dashed/solid edges are supports/atttacks. Green/red edges are supports/attacks (from arguments) with nonzero strength, and black edges are (from arguments) with zero strength. Edge width reflects the strength of originating arguments. (Strengths below/above nodes. Details in Section 6.)
@@
As the use of AI in society grows, addressing its potential unfairness is becoming increasingly crucial. For example, an unfair AI model used to support decision-making in healthcare, discriminating against a certain sub-population, would be extremely harmful. In particular, it is essential to address any potential biases against specific groups/individuals in data and AI models trained thereupon (Mehrabi et al. 2022; Caton and Haas 2024). Various notions of fairness have been proposed in the literature (Mehrabi et al. 2022; Caton and Haas 2024; Waller et al. 2024). Amongst these notions, statistical (or demographic) parity (Corbett-Davies et al. 2017) defines fairness as an equal probability of being classified with the desirable (positive) label in different groups. However, existing solutions to identify and rectify unfairness tend to focus on optimising fairness and ignore the need for transparency for the roots of (un)fairness (Caton and Haas 2024). Instead, transparency, afforded by interpretability and explainability, is crucial for the detection and mitigation of bias, as is the case for any algorithmic solution, but even more so when bias is concerned, given that fairness is inherently human-oriented (Waller et al. 2024). While considerable efforts have been made towards explainability of machine learning models (Dejl et al. 2025; Ayoobi et al. 2019, 2021, 2022, 2023a, 2023b, 2025; Leofante et al. 2024), algorithmic fairness has received limited attention (exceptions include (Grabowicz, Perello, and Mishra 2022; Waller, Rodrigues, and Cocarascu 2024)).
@@
We propose a novel transparent method to detect bias through argumentative debates about the presence of bias against individuals, based on the values of protected features for the individuals and others in their neighbourhoods. Debates recently emerged as a powerful mechanism to address various challenges in contemporary AI, such as AI safety (Brown-Cohen, Irving, and Piliouras 2024), and to improve LLM performance (Khan et al. 2024). In much of this literature, debates are unstructured and inform other entities that decide the debates’ outcome. Thus, while these debates provide a justification for the outcomes, these outcomes are not faithfully explained by the debates.
@@
Our novel method for bias detection is designed to guide debates about the presence of bias against individuals. To structure debates, we rely upon argument schemes (with critical questions) (Walton, Reed, and Macagno 2008; Macagno, Walton, and Reed 2017), from formal argumentation. To decide the debates’ outcomes, we compute arguments’ strength with gradual semantics (Baroni, Rago, and Toni 2018) from computational argumentation. In summary:
@@
The proofs of all results in the paper can be found in Appendix B. We will release the code on acceptance.
@@ 2 Related Work
Various methods for identifying, measuring and mitigating against bias have been proposed (Mehrabi et al. 2022; Caton and Haas 2024), but they often neglect transparency (Waller et al. 2024). One exception is (Grabowicz, Perello, and Mishra 2022), which proposed explanations to unearth and rectify bias, but relies upon feature importance, rather than debates, for transparency.
Bias can be defined in different ways, e.g. by comparing the success probability of the protected group to the one of unprotected individuals or the general population (statistical parity), or by comparing the risk of misclassification of the protected to the one of the non-protected group (predictive equality) (Corbett-Davies et al. 2017). Our approach is based on a local notion of statistical parity. That is, we identify neighbourhoods where groups are locally advantaged/disadvantaged with respect to success probability. The decision whether a bias exists is then based on the evidence provided by these neighbourhoods in a transparent way.
@@
(Walton, Reed, and Macagno 2008; Macagno, Walton, and Reed 2017) propose an argument scheme to capture arguments from bias111 //www.rationaleonline.com/map/9gy9gd/argument-from-bias, allowing to draw the conclusion that an arguer is unlikely to have taken both sides of an issue into account if the arguer is biased, subject to addressing critical questions regarding (i) evidence for the arguer’s bias and (ii) the need to take multiple sides for that issue. Our new argument scheme below is orthogonal to this and can be used in combination with it, by addressing the first critical question using properties of neighbourhoods to collect evidence for the bias of the arguer (a classifier).
@@
Several have pointed to the connection between argumentation and debate, e.g. (Panisson, McBurney, and Bordini 2021; de Tarlé, Bonzon, and Maudet 2022) in general, and (Rago et al. 2025) for product recommendation. While we can leverage on these works towards debate generation, our approach is orthogonal.
@@ 3 Preliminaries
For $\mathcal{D}\!=\!\bigtimes_{i=1}^{k}D_{i}$, vectors $\mathbf{x}\in\mathcal{D}$ are the inputs of the classification problem. We consider (binary) classifiers $\operatorname{c}:\mathcal{D}\rightarrow\mathcal{C}\!=\!\{0,1\}$. We regard $1$ as a desirable outcome (e.g., acceptance of a loan application) and 0 as a negative outcome.
@@
QBAFs are often compared based on properties (Leite and Martins 2011; Amgoud and Ben-Naim 2016, 2017; Baroni, Rago, and Toni 2018), many of which are tied to properties of aggregation and influence functions (Mossakowski and Neuhaus 2018). We will use properties of (strict) monotonicity and balance of these functions from (Potyka and Booth 2024b) and recap them in Appendix A. As shown in (Potyka and Booth 2024a), product and sum aggregation satisfy balance and monotonicity, sum satisfies strict monotonicity; the influence functions of DF-QuAD and quadratic energy satisfy all properties.
@@ 4 Neighbourhoods and their Properties
The goal of this section is to make the idea of local bias precise. Intuitively, a local bias against a reference individual exists if other similar individuals are advantaged by the classifier. Formally, these similar individuals can be given by a neighbourhood of individuals surrounding the reference one.
@@
Without further restrictions, an adversarial agent could define a neighbourhood by picking individuals systematically to demonstrate the (non-)existence of a bias. To prevent this, one reasonable assumption is that when we pick two individuals from a sample, we cannot leave out individuals between them. Mathematically, betweenness can be described by convexity:
@@
The next proposition explains that $S$-objectivity is always satisfied when we use an $\epsilon$-neighbourhood $\mathcal{N}_{\mathbf{x}}=\{\mathbf{x}^{\prime}\in S\ \mid\ \|\mathbf{x}-\mathbf{x}^{\prime}\|\leq\epsilon\}$ of a point $\mathbf{x}$, that can be defined w.r.t. any common distance measure, including Euclidean, Manhattan and the Hamming distance and their weighted variants.
@@
Ideally, our neighbourhoods contain good representations of both the protected and non-protected groups. If it contained only one protected individual, one may wonder about other protected individuals. For example, even for a strongly biased model, we may be able to find one protected individual that is treated similar to non-protected ones even though the model strongly discriminates against the non-protected individuals in a majority of cases. Similarly, if it contained only one non-protected individual, this may simply be an outlier. Since protected groups are often underrepresented, we should not expect a balanced neighbourhood. However, if almost all elements in a neighbourhood belong to one group, the neighbourhood may not provide very strong evidence. To quantify neighbourhood diversity, we can use entropy:
@@
The entropy $H_{\mathcal{N}}(X_{p})$ of the marginal distribution of the protected feature is defined as
$H_{\mathcal{N}}(X_{p})=-P_{\mathcal{N}}(X_{p}=g)\cdot\log P_{\mathcal{N}}(X_{p}=g)$
$-P_{\mathcal{N}}(X_{p}\neq g)\cdot\log P_{\mathcal{N}}(X_{p}\neq g).$
@@
5 Argument scheme for detecting local bias
Our method makes use of properties of neighbourhoods to drive argumentative debates about the presence of bias against $X_{p}=g$. To shape these debates, we define novel argument schemes with accompanying critical questions. In the next section, we will turn both into QBAFs.
Figure 2 gives our argument scheme for detecting local bias against the protected feature taking the value of interest ($X_{p}=g$). The critical questions address the overall quality (properties) of the neighbourhood: CQ1. Is $\mathcal{N}$ NN-significant? CQ2. Is $\mathcal{N}$ $S$-objective? CQ3. Is $\mathcal{N}$ $\epsilon$-diverse?
When the critical questions are instantiated they give attacks against the conclusion of the (instantiated) argument scheme.
@@ 5 Argument scheme for detecting local bias
We can also combine arguments drawn from different neighbourhoods, via the argument scheme in Figure 3. For simplicity, we will omit to consider any critical questions for this argument scheme, leaving this to future work.
@@
6 Arguing about Bias
Our goal is to empower debates drawn from evidence for bias. To do so, first, to identify bias in a given neighbourhood $\mathcal{N}$, we introduce our core QBAF $Q_{\mathcal{N}}$ based on the argument scheme in Figure 2. Then, we extend $Q_{\mathcal{N}}$ with critical questions. Finally, we build our QBAF $\mathcal{Q}$ that allows to argue across neighbourhoods, based on the argument scheme from Figure 3. In this section, we assume a modular gradual semantics $\sigma$ with aggregation function $\operatorname{agg}$ and influence function $\operatorname{infl}$. We also omit the QBAF subscript when applying $\sigma$.
@@ 6 Arguing about Bias
The conclusion of the argument scheme in Figure 2 amounts to two arguments, $\mathrm{Disadv}_{g}$ and $\mathrm{Adv}_{g}$, that express that individuals with $X_{p}=g$ are, resp., disadvantaged and advantaged. Their base score is set to 0. The minor premises give rise to two further arguments $\mathrm{Pos}_{g}$ and $\mathrm{Pos}_{\neq g}$. Their base score corresponds to the group’s local success probability in the neighbourhood. The major premise informs about the connection between these arguments: $\mathrm{Disadv}_{g}$ is supported by $\mathrm{Pos}_{g}$ and attacked by $\mathrm{Pos}_{\neq g}$; $\mathrm{Adv}_{g}$ is attacked by $\mathrm{Pos}_{g}$ and supported by $\mathrm{Pos}_{\neq g}$. Formally:
@@
$\mathcal{A}=\{\mathrm{Disadv}_{g},\mathrm{Adv}_{g},\mathrm{Pos}_{g},\mathrm{Pos}_{\neq g}\}$,
$\mathcal{R}^{-}=\{(\mathrm{Pos}_{g},\mathrm{Disadv}_{g}),(\mathrm{Pos}_{\neq g},\mathrm{Adv}_{g})\}$,
$\mathcal{R}^{+}=\{(\mathrm{Pos}_{g},\mathrm{Adv}_{g}),(\mathrm{Pos}_{\neq g},\mathrm{Disadv}_{g})\}$, and
$\tau(a)\!=\!\begin{cases}P_{\mathcal{N}}(c(\mathbf{X})=1\mid X_{p}\neq g),&a=\mathrm{Pos}_{\neq g}\\ P_{\mathcal{N}}(c(\mathbf{X})=1\mid X_{p}=g),&a=\mathrm{Pos}_{g}.\\ 0,&else.\end{cases}$
Figure 1 shows this QBAF (yellow, right box). Intuitively, the base score of $\mathrm{Disadv}_{g}$$\mathrm{Adv}_{g}$ is 0 as our default assumption is that individuals with $X_{p}\!=\!g$ are not disadvantaged/advantaged. The base scores of $\mathrm{Pos}_{g}$$\mathrm{Pos}_{\neq g}$ are the local success probabilities of the groups $X_{p}\!=\!g$/ $X_{p}\!\neq\!g$, and 0 if it is never the case, $1$ if it is always the case, and otherwise a value between, corresponding to the relative frequency in $\mathcal{N}$.
We show that our choice of QBAF is sensible when choosing the semantics appropriately. First, intuitively, if $P_{\mathcal{N}}(c(\mathbf{X})\!=\!1\!\mid\!X_{p}\!\neq\!g)$ and $P_{\mathcal{N}}(c(\mathbf{X})\!=\!1\!\mid\!X_{p}\!=\!g)$ are equal, then there is neither reason to accept that individuals with $X_{p}=g$ are disadvantaged nor that they are advantaged. However, as $P_{\mathcal{N}}(c(\mathbf{X})=1\mid X_{p}\neq g)$ becomes larger than $P_{\mathcal{N}}(c(\mathbf{X})=1\mid X_{p}=g)$, we should gradually accept $\mathrm{Disadv}_{g}$. Conversely, as $P_{\mathcal{N}}(c(\mathbf{X})=1\mid X_{p}\neq g)$ becomes smaller than $P_{\mathcal{N}}(c(\mathbf{X})=1\mid X_{p}=g)$, we should gradually accept $\mathrm{Adv}_{g}$. We can guarantee this behaviour by choosing aggregation and influence functions with certain properties.
@@
    monotonicity, then $\sigma(\mathrm{Adv}_{g})\!\!=\!\!0$, $\sigma(\mathrm{Disadv}_{g})\!\!=\!\!0$ or both,
@@
Second, we should expect that, as the inequality between the protected and non-protected groups increases, the strength of $\mathrm{Disadv}_{g}$$\mathrm{Adv}_{g}$ should change accordingly. To study this situation, we compare the strength values in two independent local bias-QBAFs, connecting again properties of aggregation and influence functions with the expected behaviour:
@@
Third, by construction, we treat evidence for being advantaged and disadvantaged equally. That is, in two neighbourhoods with symmetric evidence, the strength of $\mathrm{Adv}_{g}$ in one will be the strength of $\mathrm{Disadv}_{g}$ in the other:
@@
What if we detected bias only because the neighbourhood was chosen poorly? Here, we address this issue with critical questions from Section 5. We instantiate these questions by means of arguments and attacks added to local bias-QBAFs. We introduce three novel arguments, one per question: ss(ignificance) for CQ1, oo(bjectivity) for CQ2 and dd(iversity) for CQ3. By $CQ=\{s,o,d\}$, we denote the set of our novel arguments.
@@
The new arguments are unattacked and unsupported by other arguments, and thus their strength (for any $\sigma$) will be their base score. So, for $s$, the base score should be monotonically decreasing w.r.t. the size of $\mathcal{N}$ and eventually reach 0, so that small neighbourhoods provide strong attacks. For the $o$, since objectivity is a binary criterion, an indicator function is most natural. For $d$, the attack strength should be maximal when the population consists of only one group, and approach 0 as the distribution becomes uniform: this behaviour can be captured by entropy, as given in Section 4. However, since we cannot expect a uniform distribution for underrepresented groups, we should be able to rescale the entropy: our base score function for $d$ should thus be monotonically decreasing. Formally, critical questions lead to the QBAF below:

### v2

@@ Abstract
As the use of AI in society grows, addressing emerging biases is essential to prevent systematic discrimination. Several bias detection methods have been proposed, but, with few exceptions, these tend to ignore transparency. Instead, interpretability and explainability are core requirements for algorithmic fairness, even more so than for other algorithmic solutions, given the human-oriented nature of fairness. We present ABIDE (Argumentative BIas detection by DEbate), a novel framework that structures bias detection transparently as debate, guided by an underlying argument graph as understood in (formal and computational) argumentation. The arguments are about the success chances of groups in local neighbourhoods and the significance of these neighbourhoods. We evaluate ABIDE experimentally and demonstrate its strengths in performance against an argumentative baseline.
@@ 1 Introduction
As the use of AI in society grows, addressing its potential unfairness is becoming increasingly crucial. For example, an unfair AI model that supports decision-making in healthcare, discriminating against a certain sub-population, would be extremely harmful. In particular, it is essential to address any potential biases against specific groups/individuals in data and AI models trained thereupon (Mehrabi et al. 2022; Caton and Haas 2024). Various notions of fairness have been proposed in the literature (Mehrabi et al. 2022; Caton and Haas 2024; Waller et al. 2024). Amongst these notions, statistical (or demographic) parity (Corbett-Davies et al. 2017) defines fairness as an equal probability of being classified with the desirable (positive) label in different groups.
Existing solutions to identify unfairness tend to focus on optimising fairness and ignore the need for transparency for the roots of (un)fairness (Caton and Haas 2024). However, transparency, afforded by interpretability and explainability, is crucial for the trustworthy detection and mitigation of bias. While considerable efforts have been made towards explainability of machine learning models, e.g. as in (Ayoobi et al. 2023b; Dejl et al. 2025b), algorithmic fairness has received limited attention (exceptions include (Grabowicz, Perello, and Mishra 2022; Waller, Rodrigues, and Cocarascu 2024)).
@@
We propose ABIDE (Argumentative Bias Detection), a novel transparent method to detect bias through argumentative debates about the presence of bias against individuals, based on the values of protected features for the individuals and others in their neighbourhoods. Debates recently emerged as a powerful mechanism to address various challenges in contemporary AI, such as AI safety (Brown-Cohen, Irving, and Piliouras 2024), and to improve LLM performance (Khan et al. 2024). In much of this literature, debates are unstructured and inform other entities that decide the debates’ outcome. Thus, while these debates provide a justification for the outcomes, these outcomes are not faithfully explained by the debates.
Figure 1: A QBAF generated by ABIDE for the COMPAS dataset (with protected feature $X_{p}\!=\!race$ and protected group g=g\!= African-American) for two neighbourhoods of size $K=10,100$ (critical question arguments with zero strength omitted). Dashed/solid edges are supports/atttacks. Green/red edges are supports/attacks (from arguments) with nonzero strength, and black edges are (from arguments) with zero strength. Edge width reflects the strength of originating arguments. (Strengths below/above nodes. Details in Section 6.)
@@
ABIDE is designed to guide debates about the presence of bias. To structure debates, we rely upon argument schemes (with critical questions) (Walton, Reed, and Macagno 2008; Macagno, Walton, and Reed 2017), from formal argumentation. To decide the debates’ outcomes, we compute arguments’ strength with gradual semantics (Baroni, Rago, and Toni 2018) from computational argumentation. In summary:
@@
The proofs of all results and additional material can be found in the accompanying (Ayoobi et al. 2025a). The code is available at https://github.com/hamed-ayoobi/ABIDE.
@@ 2 Related Work
Various methods for identifying, measuring and mitigating bias have been proposed (Mehrabi et al. 2022; Caton and Haas 2024), but they often neglect transparency (Waller et al. 2024). One exception is (Grabowicz, Perello, and Mishra 2022), which uses explanations to unearth and rectify bias, but relies on feature importance, rather than debates, for transparency.
Bias can be defined in different ways, e.g. by comparing the success probability of the protected group to the one of unprotected individuals or the general population (statistical parity), or by comparing the risk of misclassification of the protected to the one of the non-protected group (predictive equality) (Corbett-Davies et al. 2017). ABIDE is based on a local notion of statistical parity. That is, we identify neighbourhoods where groups are locally advantaged/disadvantaged with respect to success probability. The decision whether a bias exists is then based on the evidence provided by these neighbourhoods in a transparent way.
@@
(Walton, Reed, and Macagno 2008; Macagno, Walton, and Reed 2017) propose an argument scheme to capture arguments from bias111www.rationaleonline.com/map/9gy9gd/argument-from-bias, allowing to draw the conclusion that an arguer is unlikely to have taken both sides of an issue into account if the arguer is biased, subject to addressing critical questions regarding (i) evidence for the arguer’s bias and (ii) the need to take multiple sides for that issue. Our new argument scheme below is orthogonal to this and can be used in combination with it, by addressing the first critical question using properties of neighbourhoods to collect evidence for the bias of the arguer (a classifier).
@@
Our approach follows the general idea of using argumentation towards explainability, e.g. as in (Ayoobi et al. 2019, 2021, 2022, 2023a, 2025b; Dejl et al. 2025a). Several have pointed to the connection between argumentation and debate, e.g. (Panisson, McBurney, and Bordini 2021; de Tarlé, Bonzon, and Maudet 2022) in general, and (Rago et al. 2025) for product recommendation. While we can leverage on these works towards debate generation, our approach is orthogonal to them.
@@ 3 Preliminaries
For $\mathcal{D}\!=\!\bigtimes_{i=1}^{k}D_{i}$, vectors $\mathbf{x}\in\mathcal{D}$ are the inputs of the classification problem. We consider (binary) classifiers $\operatorname{c}:\mathcal{D}\rightarrow\mathcal{C}\!=\!\{0,1\}$, and regard $1$ as a desirable (e.g., acceptance of a loan application) and 0 as a negative outcome.
@@
QBAFs are often compared on properties (Leite and Martins 2011; Amgoud and Ben-Naim 2016, 2017; Baroni, Rago, and Toni 2018), many of which are tied to properties of aggregation and influence functions (Mossakowski and Neuhaus 2018). We will use properties of (strict) monotonicity and balance of these functions from (Potyka and Booth 2024b) and recap them in (Ayoobi et al. 2025a). As shown in (Potyka and Booth 2024a), product and sum aggregation satisfy balance and monotonicity, sum satisfies strict monotonicity; the influence functions of DF-QuAD and quadratic energy satisfy all properties.
@@ 4 Neighbourhoods and their Properties
In this section, we will formalize the idea of local bias. Intuitively, a local bias against a reference individual exists if other similar individuals are advantaged by the classifier. Formally, these similar individuals can be represented by a neighbourhood surrounding the reference individual.
@@
Without further restrictions, an adversarial agent could define a neighbourhood by picking individuals systematically to demonstrate the (non-)existence of a bias. To prevent this, one reasonable assumption is that when we pick two individuals from a sample, we cannot leave out individuals between them. Mathematically, betweenness can be described by convexity.
@@
Formally, an input $\mathbf{x}\in\mathbb{R}^{n}$ is a convex combination of $\mathbf{x}_{1},\mathbf{x}_{2}\in\mathbb{R}^{n}$ if $\mathbf{x}=\lambda\cdot\mathbf{x}_{1}+(1-\lambda)\cdot\mathbf{x}_{2}$ for some $\lambda\in[0,1]$. Discrete ordinal features such as Boolean ({0,1}) or qualitative descriptions like small (0), medium (1), large (2) can be mapped to integers to match the definition. For nominal features such as color or marital status, we demand that they are equal in $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$ (and thus in $\mathbf{x}$).
The next proposition explains that $S$-objectivity is always satisfied when we use an $\epsilon$-neighbourhood $\mathcal{N}_{\mathbf{x}}=\{\mathbf{x}^{\prime}\in S\ \mid\ \|\mathbf{x}-\mathbf{x}^{\prime}\|\leq\epsilon\}$ of a point $\mathbf{x}$. It can be defined w.r.t. any common distance measuresuch as Euclidean, Manhattan and the Hamming distance and their weighted variants.
@@
Ideally, our neighbourhoods contain good representations of both the protected and non-protected groups. For example, even for a strongly biased model, we may be able to find one protected individual that is treated similar to non-protected ones even though the model strongly discriminates against the non-protected individuals in most cases. Similarly, if it contained only one non-protected individual, this may simply be an outlier. Since protected groups are often underrepresented, we should not expect balanced neighbourhoods. However, if almost all elements in a neighbourhood belong to one group, the neighbourhood may not provide very strong evidence. To quantify neighbourhood diversity, we use entropy.
@@
The entropy $H_{\mathcal{N}}(X_{p})$ of the protected feature $X_{p}$ in $\mathcal{N}$ is $H_{\mathcal{N}}(X_{p})\!=\!-P_{\mathcal{N}}(X_{p}\!=\!g)\cdot\log P_{\mathcal{N}}(X_{p}\!=\!g)-P_{\mathcal{N}}(X_{p}\neq g)\cdot\log P_{\mathcal{N}}(X_{p}\neq g).$
@@
5 Argument Schemes for Detecting Local Bias
ABIDE makes use of properties of neighbourhoods to drive argumentative debates about the presence of bias against individuals with $X_{p}=g$. To shape these debates, we define novel argument schemes with accompanying critical questions. In the next section, we will turn both into QBAFs.
Figure 2 gives our argument scheme for detecting local bias against the protected feature taking the value of interest ($X_{p}=g$). The critical questions address the overall quality (properties) of the neighbourhood: CQ1. Is $\mathcal{N}$ NN-significant? CQ2. Is $\mathcal{N}$ $S$-objective? CQ3. Is $\mathcal{N}$ $\epsilon$-diverse? When the critical questions are instantiated they give attacks against the conclusion of the (instantiated) argument scheme.
@@ 5 Argument Schemes for Detecting Local Bias
We can also combine arguments drawn from different neighbourhoods, via the argument scheme in Figure 3. For brevity, we omit to consider critical questions for this scheme.
@@
6 Arguing about Bias with ABIDE
ABIDE empowers debates on bias, in three steps.
  1. 1.
    Build a local bias-QBAF $Q_{\mathcal{N}}$ (based on Figure 2).
  2. 2.
    Add critical questions (CQ1-CQ3).
  3. 3.
    Build a global bias-QBAF $\mathcal{Q}_{\mathbf{g}}$ (based on Figure 3).
We assume a fixed modular gradual semantics $\sigma$ with aggregation function $\operatorname{agg}$ and influence function $\operatorname{infl}$ and omit the QBAF subscript when applying $\sigma$.
@@ 6 Arguing about Bias with ABIDE
The conclusion of the argument scheme in Figure 2 amounts to two arguments, $\mathrm{Disadv}_{g}$$\mathrm{Adv}_{g}$, that express that individuals with $X_{p}=g$ are, resp., disadvantaged/advantaged. Their base score is set to 0. The minor premises give rise to two further arguments $\mathrm{Pos}_{=g}$ and $\mathrm{Pos}_{\neq g}$. Their base score corresponds to the group’s local success probability in the neighbourhood. The major premise informs about the connection between these arguments: $\mathrm{Pos}_{=g}$ attacks/supports $\mathrm{Disadv}_{g}$$\mathrm{Adv}_{g}$, while $\mathrm{Pos}_{\neq g}$ supports/attacks $\mathrm{Disadv}_{g}$$\mathrm{Adv}_{g}$.
@@
$\mathcal{A}=\{\mathrm{Disadv}_{g},\mathrm{Adv}_{g},\mathrm{Pos}_{=g},\mathrm{Pos}_{\neq g}\}$,
$\mathcal{R}^{-}=\{(\mathrm{Pos}_{=g},\mathrm{Disadv}_{g}),(\mathrm{Pos}_{\neq g},\mathrm{Adv}_{g})\}$,
$\mathcal{R}^{+}=\{(\mathrm{Pos}_{=g},\mathrm{Adv}_{g}),(\mathrm{Pos}_{\neq g},\mathrm{Disadv}_{g})\}$, and
$\tau(a)\!=\!\begin{cases}P_{\mathcal{N}}(c(\mathbf{X})=1\mid X_{p}\neq g),&a=\mathrm{Pos}_{\neq g}\\ P_{\mathcal{N}}(c(\mathbf{X})=1\mid X_{p}=g),&a=\mathrm{Pos}_{=g}.\\ 0,&else.\end{cases}$
Figure 1 shows this QBAF (yellow, right box). Intuitively, the base score of $\mathrm{Disadv}_{g}$$\mathrm{Adv}_{g}$ is 0 as our default assumption is that individuals with $X_{p}\!=\!g$ are not disadvantaged/advantaged. The base scores of $\mathrm{Pos}_{=g}$$\mathrm{Pos}_{\neq g}$ are the local success probabilities of the groups $X_{p}\!=\!g$/ $X_{p}\!\neq\!g$.
We show that our choice of QBAF is sensible when choosing the semantics appropriately. If $P_{\mathcal{N}}(c(\mathbf{X})\!=\!1\!\mid\!X_{p}\!\neq\!g)$ and $P_{\mathcal{N}}(c(\mathbf{X})\!=\!1\!\mid\!X_{p}\!=\!g)$ are equal, then there is neither reason to accept that individuals with $X_{p}=g$ are disadvantaged nor that they are advantaged. However, as $P_{\mathcal{N}}(c(\mathbf{X})=1\mid X_{p}\neq g)$ becomes larger than $P_{\mathcal{N}}(c(\mathbf{X})=1\mid X_{p}=g)$, we should gradually accept $\mathrm{Disadv}_{g}$. Conversely, as $P_{\mathcal{N}}(c(\mathbf{X})=1\mid X_{p}\neq g)$ becomes smaller than $P_{\mathcal{N}}(c(\mathbf{X})=1\mid X_{p}=g)$, we should gradually accept $\mathrm{Adv}_{g}$. We can guarantee this behaviour by choosing aggregation and influence functions with certain properties.
@@
    monotonicity, then $\sigma(\mathrm{Adv}_{g})\!\!=\!\!0$ or $\sigma(\mathrm{Disadv}_{g})\!\!=\!\!0$,
@@
As the inequality between the protected and non-protected groups increases, the strength of $\mathrm{Disadv}_{g}$$\mathrm{Adv}_{g}$ should change accordingly. To study this situation, we compare the strength values in two independent local bias-QBAFs, connecting again properties of aggregation and influence functions with the expected behaviour:
@@
By design, we treat evidence for being advantaged and disadvantaged equally. That is, in two neighbourhoods with symmetric evidence, the strength of $\mathrm{Adv}_{g}$ in one will be the strength of $\mathrm{Disadv}_{g}$ in the other:
@@
The significance of the evidence given by a neighbourhood can be debatable. To take account of this, we introduce additional arguments corresponding to the critical questions: ss(ignificance) for CQ1, oo(bjectivity) for CQ2 and dd(iversity) for CQ3. By $CQ=\{s,o,d\}$, we denote the set of our novel arguments.
@@
The new arguments are unattacked and unsupported by other arguments, and thus their strength (for any $\sigma$) will be their base score. So, for $s$, the base score should be monotonically decreasing w.r.t. the size of $\mathcal{N}$ and eventually reach 0, so that small neighbourhoods provide strong attacks. For $o$, since objectivity is a binary criterion, an indicator function is most natural. For $d$, the attack strength should be maximal when the population consists of only one group, and approach 0 as the distribution becomes uniform: this behaviour can be captured by the entropy, as given in Section 4. However, since we cannot expect a uniform distribution for underrepresented groups, we should be able to rescale the entropy: our base score function for $d$ should thus be monotonically decreasing. Formally, critical questions lead to the QBAF below: